### v1

# Robust Utility-Preserving Text Anonymization Based on Large Language Models

Abstract
Text anonymization is crucial for sharing sensitive data while maintaining privacy. Existing techniques face the emerging challenges of re-identification attack ability of Large Language Models (LLMs), which have shown advanced capability in memorizing detailed information and patterns as well as connecting disparate pieces of information. In defending against LLM-based re-identification attacks, anonymization could jeopardize the utility of the resulting anonymized data in downstream tasks—the trade-off between privacy and data utility requires deeper understanding within the context of LLMs. This paper proposes a framework composed of three LLM-based components—a privacy evaluator, a utility evaluator, and an optimization component, which work collaboratively to perform anonymization. To provide a practical model for large-scale and real-time environments, we distill the anonymization capabilities into a lightweight model using Direct Preference Optimization (DPO). Extensive experiments demonstrate that the proposed models outperform baseline models, showing robustness in reducing the risk of re-identification while preserving greater data utility in downstream tasks. [^1] Our code and dataset are available at Github.
1 Introduction
Privacy protection is a fundamental societal value, enforced through various legal frameworks, e.g., the General Data Protection Regulation (GDPR) in the European Union and the California Consumer Privacy Act (CCPA) in the United States Voigt and Von dem Bussche (2017), among others. The recent advancement in large language models (LLMs) and artificial intelligence (AI) tools present both challenges and opportunities in achieving the goal.
Text anonymization is a critical method for safeguarding private and sensitive information. However, current techniques are vulnerable to disclosure threats from increasingly sophisticated Large Language Models (LLMs). Recent studies have demonstrated that these models can re-identify private information, even from texts anonymized by advanced methods Patsakis and Lykousas (2023); Staab et al. (2024a); Nyffenegger et al. (2024).
Figure 1: Anonymization examples of the Adversarial Feedback Staab et al. (2024b) (middle box) and the proposed RUPTA (bottom box) model. The red fonts mark the personally identifiable information. We highlight entities that are critical for our downstream task: occupation classification.
The first key challenge and requirement is, therefore, defending against LLM-based re-identification attacks. In defending against these powerful models, the anonymization process may compromise the utility of the resulting anonymized data in downstream tasks Mozes and Kleinberg (2021); Patsakis and Lykousas (2023). As shown in fig. 1, while the current state-of-the-art (SoTA) method, which conducts anonymization based on iterative refining according to feedback from a simulated attacker Staab et al. (2024b), can defend against re-identification attack well, it may eliminate the information crucial for the downstream task. We believe that the trade-off between privacy and data utility requires deeper understanding within the context of LLMs, in which LLMs’ re-identification capacity challenges the existing anonymization models, while if properly utilized, LLMs can help build more capable anonymization components to mitigate the discussed adversaries.
In this paper, we introduce a novel framework named Robust Utility-Preserving Text Anonymization (RUPTA), consisting of a privacy evaluator (P-Evaluator), a utility evaluator (U-Evaluator), and an optimization component.
These components are built on LLMs, where the P-Evaluator assesses re-identification risks and provides guidance to enhance anonymization robustness against re-identification attacks, the U-Evaluator gauges downstream tasks’ performance to indicate the level of preserved utility, and the optimization component iteratively edits the text based on these evaluation results to jointly optimize both objectives until pre-defined conditions are met. As shown in fig. 1, RUPTA can ensure privacy-preserving performance comparable to the SOTA method while retaining critical information necessary for accurately classifying the text pertaining to a Tennis Player.
The anonymization models based on LLMs often rely heavily on time-consuming and resource-intensive interactions with LLMs, making these models less feasible for large-scale or real-time applications. To mitigate this problem, we distill the anonymization capabilities into a lightweight model. Our experiments show that the fine-tuned lightweight model achieves a performance comparable to GPT-4, and utilizing Direct Preference Optimization (DPO) Rafailov et al. (2023) enhances the anonymization efficacy. Our main contributions are summarized as follows:
  • To the best of our knowledge, this is the first work to simultaneously optimize privacy and utility in text anonymization using SoTA LLMs, which is crucial for real-life applications.
  • We propose a novel framework for text anonymization that is built on the powerful ability of LLMs, consisting of a privacy evaluator, utility evaluator, and optimizer component, which work jointly to perform anonymization and show superior performance over the baseline models.
  • We develop more practical methods based on DPO to distill the anonymization capabilities into lightweight models with performance comparable to the teacher models.
  • We create a new dataset using the celebrity biographies from DBpedia Dan (2019) with occupation labels, serving as a practical benchmark for evaluating the impact of anonymization methods on utility. Anonymization results from LLMs are also included to aid future text anonymization research.
2 Related Work
Text Anonymization.
The task is primarily addressed through natural language processing (NLP) and privacy-preserving data publishing (PPDP) approaches. NLP methods use sequence labeling models trained on manually annotated data to identify and remove pre-defined categories of sensitive entities, such as names and phone numbers Hathurusinghe et al. (2021); Francopoulo and Schaub (2020). Rather than masking entities according to the pre-defined categories, the PPDP-based approaches mask entities according to the disclosure risk calculated through a privacy model defined by domain experts Sánchez and Batet (2016, 2017). However, most existing studies either neglect the utility of anonymized text for downstream tasks or only evaluate it post-anonymization Yermilov et al. (2023); Staab et al. (2024b), complicating the identification of a strategy that optimally balances privacy and utility. Furthermore, commonly used datasets Lebret et al. (2016); Pilán et al. (2022) in this field often lack labels for specific downstream tasks, rendering it difficult to assess the impact of anonymization operations on them.
LLMs as the Black-box Optimizer
Optimization entails the iterative generation and evaluation of solutions to enhance a specific objective function. Leveraging their robust knowledge storage and generation capabilities, LLMs can identify optimal solutions for intricate real-world optimization problems through effective prompting without necessitating additional training Prasad et al. (2023); Zhou et al. (2023). In the context of multi-objective optimization problems (MOPs), which involve two or more conflicting objectives, current methodologies typically combine Evolutionary Algorithms with LLMs Yang and Li (2023). This approach, however, requires numerous objective evaluations, rendering it impractical for scenarios where evaluating objectives is costly. Our proposed RUPTA serves as an alternative when the preference over objectives is pre-defined.
Figure 2: The framework of our proposed RUPTA method
3 Our Approach
In this section, we present our proposed RUPTA framework, which protects the privacy of the sensitive text while maintaining its utility for analytical purposes. The overview of our framework is shown in fig. 2. Given a span of text $\boldsymbol{x}_{0}$, RUPTA iteratively refines the anonymized text to optimize the privacy and utility objectives simultaneously. At iteration $t+1$, the previously anonymized text $\boldsymbol{x}_{t}$ is input into the system, as shown in the bottom left of fig. 2. The privacy evaluator (P-Evaluator) analyzes $\boldsymbol{x}_{t}$ to determine its privacy protection level based on the ground-truth personal information $y$ and provide feedback to enhance its robustness against re-identification attacks. The utility evaluator (U-Evaluator) assesses its usefulness for the downstream tasks based on the corresponding ground-truth label $c$. Feedback from both evaluators is then used by the optimizer to refine the text using available editing operations, producing the updated text $\boldsymbol{x}_{t+1}$, as shown in the top right of fig. 2. Specific content of the involved instructions can be found in section C.1.
3.1 Problem Formulation
The text anonymization challenge can be recast as a multi-objective optimization problem with two conflicting objectives: privacy and utility. In this context, privacy should be prioritized over utility. This hierarchy is established by ordering the objectives, transforming the problem into a lexicographic optimization issue Zykina (2004). The primary objective is to maximize the level of privacy preservation, ensuring that sensitive information is well-protected against re-identification risks. The secondary objective is to maintain as much useful information as possible in the anonymized text for analytical tasks. This lexicographic optimization problem can be formally expressed as
lex max $\text{lex max }F(\boldsymbol{x})=[f_{p}(\boldsymbol{x}),f_{u}(\boldsymbol{x})]$ (1)
[…]
where $f_{p}(\cdot)$ and $f_{u}(\cdot)$ denote the privacy and utility objective function, respectively. $\mathcal{X}_{0}$ denotes the set of all possible edits of $\boldsymbol{x}_{0}$. A solution $\boldsymbol{x}_{a}\in\mathcal{X}_{0}$ is lexicographically preferable to another solution $\boldsymbol{x}_{b}\in\mathcal{X}_{0}$, denoted as $\boldsymbol{x}_{a}\succ_{\text{lex}}\boldsymbol{x}_{b}$, if and only if $f_{p}(\boldsymbol{x}_{a})>f_{p}(\boldsymbol{x}_{b})$ or $(f_{p}(\boldsymbol{x}_{a})=f_{p}(\boldsymbol{x}_{b})\text{ and }f_{u}(\boldsymbol{x}_{a})>f_{u}(\boldsymbol{x}_{b}))$. To solve this lexicographic optimization problem, we propose RUPTA, an iterative method with LLMs to generate, evaluate, and optimize the anonymized text.
3.2 The Privacy Evaluator
Input Anonymized text $\boldsymbol{y}_{t}$, ground-truth personal information $x$, instruction $\mathbf{I}_{p}$, P-Evaluator $\mathcal{LLM}(\cdot)$
Output Privacy objective value $p_{t}$ and textual feedback $\boldsymbol{f}_{t}$
Algorithm 1 Privacy Objective Evaluation $f_{p}$
[…]
2:if $y$ in $(y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{K})$ then
3:     $p_{t}\leftarrow$ rank of $y$ in $(y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{K})$
4:     $\boldsymbol{f}_{t}\sim\mathcal{LLM}(\mathbf{I}_{pa}||\boldsymbol{x}||y)$
5:else
6:     $p_{t}\leftarrow K+1$
7:     $\boldsymbol{f}_{t}\leftarrow\emptyset$
8:end if
The role of the Privacy Evaluator (P-Evaluator) is to assess the privacy protection level of the anonymized text, ensuring that private content is adequately obscured against re-identification. Besides, it is essential to provide textual feedback to the LLM optimizer as guidance Pryzant et al. (2023). Thus, the privacy objective evaluation process $f_{p}(\cdot)$ is formally defined as
$\boldsymbol{f}_{t},p_{t}=f_{p}(\boldsymbol{x}_{t})$ (2)
where $p_{t}$ denotes the value of the privacy objective and $\boldsymbol{f}_{t}$ denotes the textual feedback. We describe the detailed process of privacy evaluation in algorithm 1.
P-Evaluator is instantiated as an LLM. Given the anonymized text $\boldsymbol{x}_{t}$, we concatenate it with the privacy inference instruction $\mathbf{I}_{p}$ as input to prompt the P-Evaluator to semantically infer the personal information as shown in line 1 of algorithm 1, where ||||| | denotes concatenation. This step generates top-$K$ inference results $[y^{\prime}_{i}]_{1}^{K}$ for the personal information. Each result is then compared with the ground-truth personal information $y$. If a match is found within these top-$K$ results, its rank is used as the scalar privacy score $p_{t}$. Further, the evaluator is prompted to provide natural language feedback $\boldsymbol{f}_{t}$ detailing the clues that led to the correct inference. Otherwise, we set the $p_{t}$ as $K+1$, representing the maximum achievable score for the privacy objective.
The scalar score $p_{t}$ quantifies the privacy risk associated with the anonymized text, while the textual feedback $\boldsymbol{f}_{t}$ offers qualitative insights, guiding the lexicographic optimizer on how to better obscure identifiable information. The value of $K$ serves as a customizable parameter that adjusts the sensitivity of the privacy evaluation, with higher values indicating a more inclusive search for potential privacy breaches, thus facilitating a manually adjustable trade-off between privacy and utility.
3.3 The Utility Evaluator
The Utility Evaluator (U-Evaluator) is used to ensure that the anonymized text retains its utility for specific analytical tasks, a critical consideration for practical applications across various domains. It analyzes the anonymized text $\boldsymbol{x}_{t}$, specifically assessing its effectiveness in supporting accurate occupation classification $c$. The formal utility objective evaluation process is defined as
$u_{t}=f_{u}(\boldsymbol{x}_{t},c)$ (3)
where $u_{t}$ is the utility objective value.
In this paper, we instantiate the U-evaluator with an LLM. Given the anonymized text $\boldsymbol{x}_{t}$ and the corresponding ground-truth occupation label $c$, the LLM-based U-evaluator follows the instruction $\mathbf{I}_{u}$ to output a confidence score $u_{t}$
$u_{t}\sim\mathcal{LLM}(\mathbf{I}_{u}||\boldsymbol{x}_{t}||c),$ (4)
this confidence score quantifies the evaluator’s uncertainty that $\boldsymbol{x}_{t}$ can be correctly classified into the ground truth occupation category $c_{t}$, reflecting the degree to which key utility information is preserved.
To better align feedback with real-world use scenarios, the U-Evaluator can be instantiated with the actual model employed in the downstream task. For example, if the anonymized text is intended for sentiment analysis, the U-Evaluator can be instantiated with a sentiment analysis model. The utility score $u_{t}$ can then be calculated through the logit of the ground-truth label following the traditional uncertainty quantification method Sensoy et al. (2021).
3.4 Lexicographic Optimizer
Lexicographic optimization (LO) is a special case of MOPs where multiple conflicting objectives are to be maximized simultaneously. In LO, objectives are ranked in order of importance, enabling prioritization of the most critical objectives. The LO problem is generally solved by the sequential optimization method Zykina (2004); Zhang et al. (2022). Specifically, regarding the text anonymization problem, privacy and utility are the two objectives, and privacy should be prioritized.
RUPTA employs the LLM as a black-box lexicographic optimizer in a zero-shot manner, where the LLM is prompted to achieve better solutions incrementally based on the history of optimization results and objective evaluations. The overall prompt consists of the pre-defined overall optimization description prompt $\mathbf{I}_{r}$, the memory module $\mathcal{M}$, the meta instruction variable $\boldsymbol{I}_{me}$ and the textual feedback $\boldsymbol{f}_{t}$ from P-Evaluator. The memory module $\mathcal{M}$ stores history optimization results and their corresponding privacy and utility objective values. Formally, we have =$\mathcal{M}=\{(\boldsymbol{x}_{i},p_{i},u_{i},r_{i})|i=1,2,...,t\}$.
To ensure that the primary goal of achieving maximum privacy is prioritized and only after the privacy objective is satisfactorily met does the optimizer focus on improving utility, the lexicographic-optimizer LLM operates in two different modes. When the privacy objective value has not yet reached the pre-set maximum $K+1$, the lexicographic optimizer should focus on maximizing the privacy objective, which is achieved by taking the value of meta instruction variable as $\mathbf{I}_{pr}$ that instructs the LLM to further anonymize $\boldsymbol{x}_{t}$ according to the textual feedback $\boldsymbol{f}_{t}$. The process can be formulated as
$\boldsymbol{x}_{t+1}\sim\mathcal{LLM}(\mathbf{I}_{r}||\mathcal{M}||\mathbf{I}_{pr}||\boldsymbol{f}_{t})$ (5)
Once the privacy objective value has reached the maximum threshold, the meta instruction shifts to $\mathbf{I}_{ur}$, prompting the LLM to optimize the utility level without compromising the achieved privacy objective value.
$\boldsymbol{x}_{t+1}\sim\mathcal{LLM}(\mathbf{I}_{r}||\mathcal{M}||\mathbf{I}_{ur})$ (6)
This iterative process continues until either the pre-defined maximum values for both objectives are reached or the maximum number of iterations $T$ is met. By continuously refining and evaluating the anonymized text, the optimizer iteratively improves it to achieve an optimal balance between privacy and utility.
3.5 Distilling the Anonymization Ability
Utilizing LLMs for text anonymization is computationally expensive, and for certain LLMs, access is only available through APIs, which raises privacy and cost concerns. However, in our framework, the optimization result heavily depends on the reasoning ability of the LLM, which stems from the large scale of parameters these models possess. Recent studies have demonstrated that prompting LLMs as optimizers is less effective with smaller-scale models Zhang et al. (2024).
To address this issue, we employ knowledge distillation (KD), where a large model (the teacher) transfers its knowledge to a smaller model (the student). Typically, KD involves training the student model using the outputs of the teacher model as labels Kim and Rush (2016). In our case, we utilize the final anonymization result produced by the teacher model during the lexicographic optimization as the training label for the student model.
To utilize the generation results of the teacher model more efficiently, we adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) method. This method fine-tunes an LLM on human labels of the relative quality of model generations to align the model with human preferences. In our method, intermediate optimization results from the teacher model can be considered less preferred than the final optimization result. These intermediate and final results form the preference dataset. We fine-tune the student model using DPO on this dataset to preferentially generate outputs similar to the final optimization result while reducing the likelihood of producing results akin to the intermediate stages.
4 Experimental Set-up
Dataset #Train #Validation #Test
DBPedia Classes 1938 243 239
Personal Reddit 318 - 207
Table 1: Statistics of experiment datasets.
Method Disclosure Risk Utility Preserving
[…]
DB-bio Original 100.00 98.45 99.58 99.68 99.61 99.58 0.0422
Azure Aahill (2023) 78.24 80.87 91.63 95.04 92.39 92.47 0.3202
DEID-GPT Liu et al. (2023) 77.10 79.47 90.82 94.37 92.56 91.22 0.3103
SD Dou et al. (2023) 73.21 73.63 92.27 93.11 92.69 92.96 0.2719
AF Staab et al. (2024b) 52.91 50.84 91.20 94.26 91.75 92.02 0.4048
RUPTA (Mixtral 8$\times$22b) 67.78 67.15 96.18 97.13 96.30 96.23 0.2167
RUPTA (Llama-3-70b) 64.02 63.23 95.34 96.23 95.55 95.82 0.2224
RUPTA (GPT-3.5) 68.51 69.16 95.40 96.02 95.70 95.49 0.2188
RUPTA (GPT-4) 52.67 53.11 95.58 96.26 95.91 96.02 0.1618
Table 2: Main experiment results on the test set of DB-bio dataset. The top and second performance are highlighted with bold font and underline, respectively.
Datasets.
We evaluate our text anonymization method on the following two datasets:
  • We sampled celebrity biographies from the DBpedia Classes dataset Dan (2019) to build a new dataset DB-bio. Unlike the commonly-used Wiki-bio dataset Lebret et al. (2016) in anonymization studies that lacks annotations for downstream tasks, this dataset includes detailed three-level hierarchical category annotations. We use the third-level category labels as occupation classification labels to assess the impact of our anonymization method on this specific downstream task. The name of the person described by the biography is used as the ground-truth personal information.
  • To further validate the generality of our method, we evaluate it on the PersonalReddit (PR) dataset  Staab et al. (2024a) consisting of 525 human-verified synthetic public Reddit comments and corresponding user profiles. We use the annotated occupation attribute in the profile as the label of the occupation classification task and anonymize the comments to prevent the identification of other personal attributes.
General statistics of these datasets can be seen in table 1. Detailed statistics, including category distributions, are provided in appendix A.
Evaluation Metrics.
To evaluate our text anonymization method, we focus on two critical aspects: disclosure risk and utility preservation. Disclosure risk is assessed by measuring the Success Rate (SR) of a strong adversarial LLM in inferring personal information from anonymized text. Additionally, we prompted an LLM to generate the Confidence Scores (CS), evaluating the degree of confidence with which anonymized text can be linked to the ground-truth personal information.
Utility preservation metrics are gauged by the performance of a simple neural network classifier trained on non-anonymized train data and tested on anonymized text, including Accuracy, Precision, Recall, F1 Score, and the classifier’s loss function value indicating classification uncertainty. Specific metric settings can be seen in appendix B.
Comparison Methods.
To establish the effectiveness of our text anonymization framework, we benchmark it against state-of-the-art methods and industry standards.
  • We use Azure Aahill (2023)’s industry-standard state-of-the-art text anonymizer as a traditional anonymization baseline.
  • AF Staab et al. (2024b) is a current state-of-the-art method for text anonymization based on the adversarial feedback mechanism.
  • DEID-GPT Liu et al. (2023) prompts the LLM to mask out all the entities of pre-defined kinds.
  • SD Dou et al. (2023) prompts the LLM to replace the entities of pre-defined kinds with more general counterparts.
All these methods are recreated using the GPT-4 model Achiam et al. (2023). Besides, we explore the effectiveness of using different LLM architectures as the lexicographic optimizer, including open-sourced models like instruction-tuned Llama-3-70b AI@Meta (2024) and Mixtral $8\times 22$Jiang et al. (2024), and the proprietary GPT-4 and GPT-3.5. Besides, we evaluate the original non-anonymized dataset (Original) for reference.
Implementation Details.
GPT-4 is used exclusively as the privacy evaluator of RUPTA and simulated attacker of AF due to its advanced capabilities in re-identification. GPT-4 is also used as the utility evaluator of RUPTA. Besides, we experimented with using Phi-3 Mini Abdin et al. (2024) and Llama-3-8b AI@Meta (2024) as the student model. Details can be seen in appendix C.
5 Experimental Results
5.1 Overall Results
Method Disclosure Risk Utility Preserving
[…]
Personal Reddit Original 49.76 81.89 55.13 63.51 55.80 58.45 1.5695
Azure Aahill (2023) 45.89 81.07 54.04 58.49 54.17 57.00 1.7340
DEID-GPT Liu et al. (2023) 43.12 72.81 53.98 58.21 54.06 56.31 1.9314
SD Dou et al. (2023) 44.05 75.17 54.11 58.43 54.21 56.93 1.7501
AF Staab et al. (2024b) 35.40 57.76 16.64 22.32 16.68 21.26 3.3380
RUPTA (Mixtral 8$\times$22b) 35.27 65.56 37.37 47.82 37.67 43.48 2.2836
RUPTA (Llama-3-70b) 39.61 61.63 32.96 44.57 32.82 38.65 2.3131
RUPTA (GPT-3.5) 34.30 61.50 32.04 40.44 31.97 36.23 2.4477
RUPTA (GPT-4) 35.75 55.04 30.34 39.14 30.09 35.75 2.5391
Table 3: Experimental results on the test set of PersonalReddit dataset. The top and second performance are highlighted with bold font and underline, respectively.
The overall experimental results on the DB-bio dataset are presented in table 2. In the disclosure risk evaluation, methods that anonymize the data in an iterative refinement manner, including our RUPTA method and the AF method, achieve the best performance. Although DEID-GPT and SD also leverage LLMs, they follow a traditional approach focusing on masking entities of pre-defined types. Experiment results demonstrate that such methods cannot adequately defend against re-identification attacks from LLMs. Additionally, using open-source LLMs as the lexicographic optimizer also achieves comparable privacy-preserving performance, demonstrating the practicality and generality of our method.
For the utility preserving evaluation, traditional methods like Azure mask all the entities of pre-defined kinds with “*”, leading to the most significant information loss, thus achieving the lowest performance. The DEID-GPT, SD, and AF methods, although anonymized through replacing sensitive entities with more general ones, do not consider the downstream analysis task and generalize all the possible sensitive entities, which also significantly undermines downstream task performance. Visualization results of the optimization process in fig. 3 highlight the drawback of the AF method, where the SR and classification accuracy decrease simultaneously as the number of optimization steps increases. In contrast, our method achieves the best downstream task performance. Furthmore, during the optimization process of RUPTA, there is an explicit increasing phase of the classification accuracy, demonstrating the effectiveness of the RUPTA method to maximize both the privacy and utility in the anonymization process. This trend also illustrates that beyond a certain point, further anonymization yields diminishing returns in privacy preservation and results in greater losses of utility information.
Figure 3: Evaluation results of the anonymized text at each iteration during the anonymization process using the AF and RUPTA methods with GPT-4, Llama-3-70b (Llama-3), and Mixtral $8\times 22$b (Mixtral) as optimizers on the test set of the DB-bio dataset.
5.2 Customizable Privacy-Utility Tradeoff
Figure 4: Customizable privacy-utility tradeoff experiments on the test set of DB-bio dataset with GPT-4, Llama-3-70b (Llama-3), and Mixtral $8\times 22$b (Mixtral) as optimizers, respectively.
The experiment results for the customizable privacy-utility tradeoff are displayed in fig. 4. In our method, the maximum value of the privacy objective is set manually according to specific requirements, allowing for a customizable privacy-utility tradeoff. We analyze and visualize the average SR and classification accuracy of our method using GPT-4, Llama-3-70b, and Mixtral $8\times 22$b as the lexicographic optimizer. We set the maximum privacy value to 1, 5, 10, 15, and 20, respectively. It is evident in fig. 4 that our proposed method can effectively adapt the privacy-preserving level according to the maximum value setting. As the maximum privacy value increases, the average privacy score improves while the utility score adjusts accordingly. This observation demonstrates the flexibility of our approach in balancing privacy and utility based on user-defined requirements.
5.3 Experiments on PR Dataset
To evaluate the generality of our method, we further conduct experiments on the PR dataset with results presented in table 3. The PR dataset is characterized by fewer explicit and more implicit sensitive entities. Entity recognition-based methods, including Azure, DEID-GPT, and SD, struggle to detect these implicit entities, resulting in minimal masking operations, as evidenced by their evaluation results closely mirroring those of the original dataset. Consequently, while these methods exhibit higher performance on the downstream task, they provide inferior privacy protection. Only the AF and our method can properly detect implicit sensitive information and achieve the lowest disclosure risk. However, the AF method anonymizes without tailoring its approach to the specific downstream task, which significantly impairs task performance. In contrast, our method not only effectively minimizes disclosure risk but also preserves a greater degree of utility in the anonymized text than AF, achieving a better privacy-utility tradeoff.
5.4 Distilled Models
In this experiment, we try to distill the anonymization ability of GPT-4 into lightweight models. Using RUPTA with GPT-4 as the lexicographic optimizer, we anonymized the training and validation sets of the DB-bio dataset. Initially, we fine-tuned student models in a supervised manner (SFT) using the final optimization results as labels. Then we constructed a preference dataset from the optimization trajectories and conducted DPO fine-tuning on the optimal checkpoint during the SFT phase. The evaluation results are presented in fig. 5.
Figure 5: Knowledge distillation experiment results using Llama-3-8b (Llama-3) and Phi-3 Mini (Phi-3) as the student model, respectively.
From the disclosure risk evaluation, we observe that the primarily supervised fine-tuning on the final optimization results enables the smaller models to achieve performance comparable to the teacher model, GPT-4. Additionally, the DPO fine-tuning process further enhances the performance of the student models, narrowing the gap to the teacher model’s capabilities.
In the utility preserving evaluation results, in addition to the classification accuracy, we further demonstrate the semantic similarity between the anonymized and original text. The supervised fine-tuned student models maintain a high level of downstream task performance. Although the DPO fine-tuning process improves the privacy-preserving performance, it somewhat harms the downstream task performance. This phenomenon likely results from the unbalanced optimization phases in the lexicographic optimization process, where achieving the maximum privacy objective requires more iterations than improving downstream task performance, as shown in fig. 3. Consequently, the student models, fine-tuned with DPO, prioritize privacy to a greater extent, potentially at the expense of utility. Anonymization examples are shown in fig. 6. We can see that the student model can learn to generalize or remove sensitive entities after the SFT phase. After the DPO fine-tuning phase, the student model can further generalize sensitive entities marked by underlining, e.g., from “father” to “family member”. Both models can keep the relevant information about the downstream task in the anonymized text, as highlighted in the figure.
Figure 6: Anonymization example of Phi-3 Mini model
6 Conclusions
This paper presents a novel framework that integrates a privacy evaluator, a utility evaluator, and an optimizer to effectively anonymize text for text anonymization using LLMs, ensuring reduced risk of re-identification while maintaining utility for downstream tasks. Building on that, we further develop practical methods based on DPO to distill the anonymization capabilities into lightweight models with the performance comparable to that of the teacher models. Additionally, the creation of a new dataset based on celebrity biographies with occupation labels provides a valuable resource for assessing the impact of various anonymization techniques on the specific downstream task-occupation classification. The superiority of our methods over existing models contributes to text anonymization and sets new baselines for future research that considers downstream utility in anonymization.
Limitations
While our study presents significant advancements in text anonymization techniques using LLMs, there are several limitations to acknowledge and to be mitigated in the future work.
Firstly, the reliance on LLMs, while beneficial for capturing complex patterns and associations, also makes our approach computationally intensive, potentially limiting its applicability in environments with constrained computational resources, despite the use of a distilled, lightweight model.
Secondly, our framework’s performance, though superior to baseline models, still depends heavily on the quality and diversity of the training data. The new dataset derived from celebrity biographies may not fully represent the variety of scenarios in which text anonymization is needed, potentially affecting the generalizability of our findings to other domains or more diverse datasets.
Besides, our approach assumes a static adversarial model where the capabilities of potential adversaries are constant. However, in real-world scenarios, adversaries may evolve, adopting more sophisticated techniques to re-identify data. This dynamic aspect of threat models poses a significant challenge, as our framework might not fully account for the adaptive strategies of adversaries over time. To address this, continuous updates and iterative improvements to the framework will be necessary to maintain robustness against emerging re-identification methods.
Lastly, a critical limitation of our method, as well as all NLP-based anonymization approaches, is the absence of formal guarantees of the privacy protection level. While traditional Named Entity Recognition (NER)-based methods struggle with the nuanced capabilities of modern LLMs, our approach, and similarly the AF method, provide an experimental metric demonstrating reduced re-identification risk when contending with state-of-the-art LLMs like GPT-4. Currently, offering a formal guarantee for NLP-based anonymization methods remains challenging; instead, providing an experimental guarantee seems more feasible. This could involve assessing to what extent an anonymization method can defend against re-identification attacks from current LLMs, which have demonstrated formidable re-identification capabilities due to their extensive knowledge stored in parameters. Future work could aim to establish a general metric for this experimental guarantee, potentially linking this risk metric with human perceptions or requirements for text quality and privacy protection levels, through methods such as conducting human evaluations. These limitations underscore the need for ongoing research to refine these approaches, enhance their adaptability, and address the broader implications of their use.
Ethics Statement
This research adheres to ethical guidelines in the development and application of text anonymization technologies using LLMs. Recognizing the dual-edged nature of anonymization—its potential to protect privacy while also possibly enabling data misuse—we have implemented several safeguards to ensure responsible use. We commit to transparency in our methodologies and the limitations of our models, as detailed in previous sections of this paper. By openly discussing the strengths and weaknesses of our approach, we aim to foster an informed community that can critically assess and improve upon our work. Besides, while developing our dataset from celebrity biographies, we have ensured that all data used were sourced from publicly available, non-sensitive information. The dataset complies with all applicable data protection laws and ethical standards, and no personally identifiable information was used without consent.
Acknowledgement
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
Appendix A Dataset Settings
Figure 7: Label distribution of the DB-bio dataset.
To build DB-bio dataset, we sampled data samples from the DBPedia Classes dataset, where each sample consists of the biography, the profile of the described people and the three-level category. We sampled according to the third level category. Specifically, we chose 24 categories, and the number of data samples for each category is shown in fig. 7. Then we manually checked each sample to filter out non-English tokens and examples with a biography longer than 700 words or shorter than 300 words. Finally, we divided the whole dataset into train, validation and test part following the ratio of 8:2:1.
Appendix B Evaluation Metrics
To evaluate our text anonymization method, we focus on two critical aspects: disclosure risk and utility preservation. Disclosure risk is assessed by measuring the success rate (SR) of a strong adversarial LLM in inferring personal information from anonymized text. A lower success rate indicates lower disclosure risk. Different from the P-Evaluator in the anonymization process, a more rigorous case is used in the evaluation set-up, where the ground-truth is mixed with other similar items and the adversarial LLM is prompted to choose one from these items according to the anonymized text. Additionally, we further prompted an LLM to generate the Confidence Scores (CS), evaluating how confidently the anonymized text can be associated with the ground-truth personal information, providing a measure of uncertainty while making inferences
Utility preservation metrics are gauged by the performance of a simple neural network classifier trained on non-anonymized train data but tested on anonymized data, including Accuracy, macro averaged Precision, macro averaged Recall, macro averaged F1 Score, and the classifier’s loss function value indicating classification uncertainty. For the DB-bio dataset, we train a BERT model Devlin et al. (2019) on the train set using validation set for hyper-parameter tuning. In the training process, we set the batch as 16 learning rate as 1e-5. We use the linear learning rate scheduler. We train the model for 20 epochs. For the PersonalReddit dataset, we train a RoBERTa-large Liu et al. (2019) model on the train set and use the test set for hyper-parameter tuning. In the training process, we set the batch as 8, learning rate as 1e-5. We use the linear learning rate scheduler. We train the model for 10 epochs.
Appendix C Implementation Details
C.1 Prompts
Figure 8: The prompt template used in the privacy evaluator to get the privacy objective value.
Figure 9: The prompt template used in the privacy evaluator to get the textual feedback.
Figure 10: The prompt template used in the utility evaluator to get the utility objective value.
Figure 11: The prompt template used in the lexicographic optimizer to optimize the anonymized text.
Figure 12: Meta instruction used in the privacy optimization phase.
Figure 13: Meta instruction used in the utility optimization phase.
Figure 14: The prompt template used to evaluate the confidence score
Figure 15: The prompt template used to generate the similar candidates used to evaluate the attack success rate.
Figure 16: The prompt template used to select from the candidate list to evaluate the attack success rate.
For the DB-bio dataset, the prompt template used in the privacy evaluator $\mathbf{I}_{p}$ is set as shown in fig. 8. The instruction used to get the textual feedback from privacy evaluator $\mathbf{I}_{pa}$ is set as shown in fig. 9. The prompt template used in the utility evaluator $\mathbf{I}_{u}$ is as shown in fig. 10. The prompt template used in the lexicographic optimizer $\mathbf{I}_{r}$ is set as shown in fig. 11. The meta instruction $\mathbf{I}_{pr}$ in the privacy optimization phase is set as shown in fig. 12. The meta instruction $\mathbf{I}_{ur}$ in the utility optimization phase is set as shown in fig. 13. The prompt template used to evaluate the confidence score metric is shown in fig. 14. The prompt template used to generate the candidate list to evaluate the success rate metric is shown in fig. 15. The prompt template used to evaluate the success rate metric is shown in fig. 16.
Figure 17: The prompt template used in the privacy evaluator to get the privacy objective value.
Figure 18: The prompt template used in the privacy evaluator to get the textual feedback.
Figure 19: The prompt template used in the utility evaluator to get the utility objective value.
Figure 20: The prompt template used in the lexicographic optimizer to optimize the anonymized text.
Figure 21: Meta instruction used in the privacy optimization phase.
Figure 22: Meta instruction used in the utility optimization phase.
Figure 23: The prompt template used to evaluate the confidence score
Figure 24: The prompt template used to generate the similar candidates used to evaluate the attack success rate.
Figure 25: The prompt template used to select from the candidate list to evaluate the attack success rate.
Figure 26: The prompt template used to choose from the pre-defined options list to evaluate the attack success rate.
For the PersonalReddit dataset, the prompt template used in the privacy evaluator $\mathbf{I}_{p}$ is set as shown in fig. 17. The instruction used to get the textual feedback from privacy evaluator $\mathbf{I}_{pa}$ is set as shown in fig. 18. The prompt template used in the utility evaluator $\mathbf{I}_{u}$ is as shown in fig. 19. The prompt template used in the lexicographic optimizer $\mathbf{I}_{r}$ is set as shown in fig. 20. The meta instruction $\mathbf{I}_{pr}$ in the privacy optimization phase is set as shown in fig. 21. The meta instruction $\mathbf{I}_{ur}$ in the utility optimization phase is set as shown in fig. 22. The prompt template used to evaluate the confidence score metric is shown in fig. 23. The prompt template used to generate the candidate list to evaluate the success rate metric is shown in fig. 24. The prompt template used to evaluate the success rate metric is shown in fig. 25. For the personal attribute with pre-defined categorical options like sex, we used the prompt template shown in fig. 26 to evaluate the success rate metric.
C.2 Knowledge Distillation
We access GPT-3.5 and GPT-4 through the API provided by Azure. We fine-tuned the two student models using the QLORA method Dettmers et al. (2024). We use the turbo version of GPT-4 for cost saving. For both the SFT and OPT fine-tuning phase, we follow the instruction fine-tuning manner where the instruction "Please anonymize the following biography:" is prepended to the input biography. For the Phi-3 Mini model, we use the released instruction-tuned version of it, we set the learning rate as 2e-4, set the batch size as 4, set the gradient accumulation steps as 4, and the epochs number as 7. The rank and alpha of the QLORA method are set as 32 and 64, respectively. The dropout rate is set as 0.05 For the Llama-3-8b model, we use the released instruction-tuned version of it, we set the learning rate as 1e-4, set the batch size as 4, set the gradient accumulation steps as 4, and the epochs number as 7. The rank and alpha of the QLORA method are set as 32 and 64, respectively. The dropout rate is set as 0.1 For both models, we quantize them with 4 bits. We use the paged adamw 32 bit optimizer and cosine learning rate scheduler. The warmup ratio is set as 0.05. The experiments are conducted on a Nvidia A100 80G GPU.
Appendix D Detailed Related Work
D.1 Text Anonymization
Text anonymization is crucial for protecting privacy in textual data, primarily addressed through natural language processing (NLP) and privacy-preserving data publishing (PPDP) approaches. NLP methods use sequence labeling models trained on manually annotated data to identify and remove pre-defined categories of sensitive information, such as names and phone numbers Hathurusinghe et al. (2021); Francopoulo and Schaub (2020); Adams et al. (2019); Eder et al. (2022); Arranz et al. (2022); Jensen et al. (2021); Kleinberg et al. (2022). NLP approaches typically do not account for non-predefined sensitive information and apply uniform masking to all detected data, lacking flexibility in adjusting the level of anonymization based on disclosure risk.
Privacy-preserving data publishing (PPDP) focuses on developing computational techniques to release data without compromising privacy. The PPDP-based approaches to anonymization is fundamentally privacy-first, enforcing a pre-defined privacy model through various data masking methods such as noise addition or value generalization Chakaravarthy et al. (2008); Cumby and Ghani (2011); Anandan et al. (2012); Sánchez and Batet (2016, 2017). For instance, the well-known k-anonymity privacy model Chakaravarthy et al. (2008) requires that each combination of quasi-identifier attribute values is shared by at least k records in the dataset. However, these methods often impractically assume that sensitive entities are pre-detected or require extensive external data resources to calculate disclosure risk Sánchez and Batet (2016), which limits their practicality in dynamic environments.
The extraordinary capabilities of LLMs significantly influence text anonymization studies. On the one hand, LLMs’ in-context learning ability have diminished the need for manually annotated training data, simplifying domain adaptation in text anonymization tasks Liu et al. (2023); Dou et al. (2023); Albanese et al. (2023). However, the powerful abilities of LLMs also introduce new threats to privacy. Their capacity to semantically infer personal information from texts provided at inference time poses a significant disclosure risk to existing anonymization techniques  Nyffenegger et al. (2024); Staab et al. (2024a); Patsakis and Lykousas (2023), which is largely overlooked both by traditional anonymization methods and emerging LLM-based approaches. In response, a concurrent study by Staab et al. introduced an Adversarial Feedback framework, where one LLM anonymizes texts based on adversarial feedback from another LLM tasked with re-identifying the text, aiming to mitigate re-identification risks from LLMs. Despite its effectiveness in enhancing privacy, this method does not account for the impact on downstream analysis, often compromising the utility of the anonymized text for further use.
D.2 Prompt Optimization with LLMs
The use of LLMs for optimization tasks has gained considerable attention, particularly in the context of prompt optimization, which refers to the process of refining the input prompts given to LLMs to maximize their performance on specific tasks. There have been many recent advancements in this area Prasad et al. (2023); Zhou et al. (2023); Xu et al. (2022); Yang et al. (2024), which have shown the potential for optimization solely through prompting without the need for additional training. While these methods achieve impressive results, they primarily focus on improving task performance without considering other important factors like instruction length and perplexity.
To address this limitation, Yang and Li formulated prompt optimization as an evolutionary multi-objective optimization problem. Using an Evolutionary Algorithm, they obtained the Pareto optimal set of prompts, allowing users to choose prompts based on their preferences over multiple criteria. Analogously, the task of text anonymization can also be framed as an multi-objective optimization problem with two conflicting objectives: privacy and utility. Different from prompt optimization, text anonymization explicitly prioritizes privacy and requires a unique optimal anonymization solution for each document. Therefore, we propose to frame text anonymization as a lexicographic optimization problem and leverage LLMs to solve it.